A conductor of length šæ moves at a steady speed vat right angles to a uniform magnetic fieldĀ of flux density šµ. Show that the e.m.f. šø across the ends of the conductor is given by theĀ 

equation šø = šµš‘™š‘£.Ā 


1
Expert's answer
2021-03-29T09:02:37-0400

Let the velocity of the rod be "v \space m/s"

Magnetic flux density = B T/m2

Length of rod = "l" m


Any change in magnetic flux induces an emf opposing that changeā€”a process known as induction. Motion is one of the major causes of induction.

Charges moving in a magnetic field experience the magnetic forceĀ "F=qvBsin\theta" , which moves opposite charges in opposite directions.


From Faraday's law,

Magnitude of emf induced along the moving rod = "N\dfrac{\Delta\Phi }{\Delta t}" (1)

As the number of turns in this coil is equal to one

"\therefore" N=1

Flux, "\Phi=BAcos\theta"

Since, magnetic field is perpendicular to the area of cross-section

"\theta=0\degree,\space cos0\degree=1"

"\Delta\Phi=\Delta (BA)\\\Rightarrow\Delta\Phi=B\Delta A\\\Rightarrow\Delta\Phi=Bl\Delta x"


Substituting the value of N and "\Delta\Phi" in equation (1),

we get,

emf "=Bl\dfrac{\Delta x}{\Delta t}"

Since "\dfrac{\Delta x}{\Delta t}=v," the velocity of the rod

"\therefore \space emf=Blv" ("B,l" and "v" are perpendicular to each other)

The above expression gives the motional emf.